OpenStudy (anonymous):

[5.02] MA.6.A.3.2 Daniel set a goal for the number of servings of fruits and vegetables he wants to eat each day. He has already eaten two and three-fourths servings today, and he still needs to eat three and three-fourths servings to meet his goal. What is his goal for daily servings of fruits and vegetables? 5 five and one-half 6 six and one-half

OpenStudy (swolesammy):

six and one half

OpenStudy (texaschic101):

2 3/4 + 3 3/4 = 11/4 + 15/4 = 26/4 = 6 1/2

OpenStudy (anonymous):

Thank U